OpenStudy (anonymous):

please help! what is the solution of the equation? -10 + square root x+8 = -4 A. 36 B. 28 C. 44

OpenStudy (shamil98):

\[-10 +\sqrt{x+8}= -4\] add 10 to both sides \[\sqrt{x+8} = 6\] square both sides\[(\sqrt{x+8})^2= 6^2\] x + 8 = 36 subtract 8 to both sides x = 28

OpenStudy (anonymous):

thank you. can you help me with other questions